OpenStudy (anonymous):

If I have a tube with ID 4” and length of 55” how can I calculate the inter surface of it assuming it is a flat sheet

OpenStudy (anonymous):

the inner surface will be the surface area of a cylinder. I believe the formula is pi*radius*hieght so the answer would be 55*4*pi == 220pi

OpenStudy (dumbcow):

Surface area of cylinder = 2*pi*r*h im assuming ID means inner diameter radius = 1/2*diameter SA = 2*2*55*pi SA = 220pi